# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 3.3 Model Inversion
Model inversion attacks generally exploit the output of the model and use this information to reconstruct the input data that generated the observed outputs. In the initial formulation (Fredrikson et al., 2014; 2015), an adversary is given access to the trained model and to a partially masked input, for which only non-sensitive features are disclosed. The attacker aims to infer the value of the sensitive attributes that serve as input to the model. While the initial formulation is implemented for linear regression models (Fredrikson et al., 2014), this type of attack has been successfully used to extract sensitive input information from different models. Fredrikson et al. (2015) applied for the first time a pattern inversion attack against a neural network, trained for facial recognition: they were able to partially reconstruct the faces of people in the training set exploiting the prediction of the system, framing the inversion as an optimization problem that requires reconstruction of the image that justifies the observed output distribution. Model inversion attacks are often performed against distributed models: in fact, some protocols like federated learning ensure that private data are not shared across multiple servers (see Section 2.2 for further details). However, with some access to model outputs, it is possible for an attacker to reconstruct the input data that causes the observed output.
@@
Model Inversion Attacks This family of attacks threatens models especially when they are distributed across multiple servers: in fact, while data holders do not directly share private data – which are thus maintained locally– they may be willing to share seemingly opaque information related to a machine learning model, such as model outputs, embeddings or gradients. However, Model Inversion attacks are designed to reconstruct private data from these apparently opaque sources.
3.3.1 Model Output Inversion
@@
The outputs of a model, like the representation of textual data that a Transformer-based LM produces in the last layer, can be totally or partially inverted to reconstruct the text that is originally the input of the model. In particular, these representations – often called embeddings – are only apparently opaque and can be inverted to reconstruct the entire sentence or the words that it contains. Pan et al. (2020) design the first model inversion attack against Transformer-based Language Models. They observe that directly framing the inversion problem as an optimization problem is challenging due to the discrete nature of sentences. They train an attack classifier that is designed to extract fixed patterns and keywords – like, for example, location or medical information – from sentence representations of models. Similarly, Song & Raghunathan (2020) predict unordered sets of words from sentence embeddings: their attack model is a neural network trained with the multiset prediction loss (Welleck et al., 2018) so that it is possible for the model to predict each word in the embedding also conditioned on the words already predicted.
@@
In some cases, the entire input text is also reconstructed. Li et al. (2023c) implement the inversion as a generative language modeling task with a decoder-only model, conditioned on the embedding of the sentence to invert as the first token representation. They show that the model successfully inverts meaningful tokens containing a large fraction of the named entity. Morris et al. (2023) propose an iterative method to reconstruct the input text of Transformer-based models that produce embeddings for documents: their attack is based on the idea of iteratively generating different hypotheses that may justify the observed embeddings. They learn the distribution of texts in a dataset given their embeddings: their approach starts by assuming an initial hypothesis and then iteratively improves it by adjusting the hypothesis to progressively align its embedding to the target one.
@@
This kind of attack has also been applied to Language Models fine-tuned on classification tasks to reconstruct private training data. In this case, the output of the model is a probability distribution over each of the possible classes. This information can be effectively exploited: in particular, an attacker may exploit the greater confidence that a model has in classifying sentences that are already part of the training set. Elmahdy et al. (2022) and Parikh et al. (2022) inject canaries in the training data (sequences that are intended to be later extracted and out of distribution with respect to the other data) and then reconstruct a partially masked sentence to search for tokens that maximize the probability of the target label. Elmahdy et al. (2022) perform the reconstruction by enumerating tokens in the vocabulary, weighted by the number of occurrences that they have in the training data. Parikh et al. (2022) propose an alternative methodology to complete the canary: they replace the discrete tokens choice with a distribution of probability – that has to be learned – over the embedding of the vocabulary for each position in the canary to be reconstructed. However, they only perform experiments against an LSTM model.
Zhang et al. (2022c) propose instead to build a dataset that resembles the unknown training data and initially train a text generator on these collected data, similarly to what is done with shadow model MIA (Section 3.2). Then, they further train the text generator via word embedding perturbation with the objective of minimizing the classification loss on the target model. The idea is that the target model’s prediction on training sentences will exhibit a lower loss and this information can be used to perturb the text generator model and obtain data closer to the private one.
@@ 3.3.2 Gradient Inversion
In federated learning (Ray et al., 2021) multiple nodes train a model cooperatively without directly sharing data samples: the system is designed to make training feasible while exchanging model updates rather than data across nodes (see §5.3 for further details). Gradient Inversion attacks are directed against federated learning environments that perform training by sharing gradients. Those attacks can be used to reconstruct training data from seemingly opaque sources.
@@
Zhu et al. (2019) demonstrated that if the gradients are openly accessible, it is possible to reconstruct the training data. The attacker aims to reconstruct the training data by minimizing the distance between some generated gradients and the ground truth one. In particular, the attacker initially generates gradients from random inputs and labels. Then, having access to the ground truth gradients solves the optimization problem of minimizing the $L_{2}$ distance between the generated gradients and the ground truth one, parameterized by real inputs and labels. In their experiments, Zhu et al. (2019) applied this attack against a BERT model. This idea was further explored by inverting gradients for image classification tasks (Zhu & Blaschko, 2020; Geiping et al., 2020; Yin et al., 2021).
@@
In reconstructing textual data, Deng et al. (2021) further attempted to perform a gradient attack against Transformer-based language models introducing also a $L_{1}$ term in the reconstruction loss. Experiments on a Transformer model, BERT model and TinyBERT demonstrate the effectiveness of this gradient attack, recovering up to 50% of the original tokens on average. In LAMP (Balunovic et al., 2022), more effective results are achieved by simultaneously training the attack model to minimize the difference between the reconstruction gradients and choosing at each iteration only sequences that have low perplexity according to an external language model (like GPT-2). Reconstructing the correct word order is, in fact, one of the main challenges: Gupta et al. (2022) propose to recover from the gradients a bag of words for the sentence to extract and then perform a beam search to effectively reconstruct the sentence.
@@ 4 Solutions for Preserving Privacy in LLMs: Data
Figure 4: Classical Anonymization Pipeline. The figure illustrates the standard process for anonymizing data in natural language processing tasks. The pipeline consists of three sequential blocks: Preprocessing, Identification, and Anonymization. The Preprocessing block prepares the raw data for further analysis. The Identification block involves the use of a procedure to detect potential Personally Identifiable Information (PII) within the data (e.g., Named Entity Recognition). Finally, the Anonymization block replaces identified PII with anonymous substitutes, ensuring privacy preservation in the data. Arrows indicate the flow of data through each stage of the process.
After identifying the key vulnerabilities and potential attacks on LLMs as outlined in §3, in this section, we introduce the reader to the current possible solutions to the problem of preserving privacy in the context of LLMs and text generative models. We have to say that the field is very fragmented as this is a very recent concern and the field is evolving fast. There is not, at the moment of writing this survey, a single gold standard solution to the problem of preserving privacy in LLMs and text generative models. The main solutions available can be broadly categorized into two categories: i) Preserving Privacy in Data §4 and ii) Preserving Privacy in Model §5. In the former section, we collected all the methods that try to solve the problem at its source i.e., they act directly on the privacy-sensitive information in the dataset, before the training happens, while in the latter we selected all the methods that address the problem from the model perspective i.e., try to remove from the model problematic information after the training. The two approaches fit different use cases, so it is important to be aware of both.
In this section, we propose several solutions that allow preserving privacy from the data perspective i.e., we preserve privacy directly in the training data used in machine learning models. This includes any approach that involves manipulating, carefully choosing, or generating training data in order to avoid the memorization and subsequent leakage of private information inside a large language model or any model that learns from data. This basically puts the focus on preserving privacy in the data used to train the model. Given that a machine learning model is moulded by its training dataset, training it on a privacy-preserving dataset ensures that these methods offer equivalent privacy assurances to the dataset itself. The section is structured in the following way: we first show classical anonymization techniques §4.1, then we focus on anonymization techniques built with Differential Privacy §5. Putting this last section in relation with Figure 6, the point of application of DP would be (1), the input, because all these techniques aim to apply Differential Privacy on the data before feeding it to the LLM (for either training or inference).

### v2

@@ 3.3 Model Inversion and Stealing
Model inversion attacks generally exploit the output of the model and use this information to reconstruct the input data that generated the observed outputs. In the initial formulation (Fredrikson et al., 2014; 2015), an adversary is given access to the trained model and to a partially masked input, for which only non-sensitive features are disclosed. The attacker aims to infer the value of the sensitive attributes that serve as input to the model: since deep learning models create dense, vectorial representations for textual data, model inversion attacks against deep learning architectures aim to decode human-readable features from them. While the initial formulation is implemented for linear regression models (Fredrikson et al., 2014) and a neural network trained for image processing (Fredrikson et al., 2015), this type of attack has been successfully used to extract sensitive input information from different models, including LSTMs (Welleck et al., 2018). Those kinds of attacks are challenging against LMs, especially since, in principle, the optimization process to reconstruct the initial input of a model has to deal with the discrete nature of text (Parikh et al., 2022).
Model inversion attacks are often performed against distributed models; in fact, some protocols like federated learning ensure that private data are not shared across multiple servers (see §2.3 for further details). However, with some access to model outputs, it is possible for an attacker to reconstruct the input data that causes the observed output.
@@
Model Inversion Attacks and Model Stealing This family of attacks threatens models, especially when they are distributed across multiple servers: in fact, while data holders do not directly share private data – which are thus maintained locally– they may be willing to share seemingly opaque information related to a machine learning model, such as model outputs, embeddings or gradients. However, Model Inversion attacks are designed to reconstruct private data from these apparently opaque sources. In some cases, also centralized private models parameters can be directly reconstructed: this is the objective of Model Stealing attacks.
In some cases, from those opaque sources, it’s also possible to reconstruct a model’s parameters from its output: in this case, the attack technique is often defined as Model Stealing (§3.3.3). These attacks aim to recover model weights from an API-gated model, making it less opaque and, hence, less robust against other types of privacy attacks. While the stealing of entire LLMs has not yet been described, some works efficiently reconstruct a subset of the model parameters in black-box attacks.
3.3.1 Model Output Inversion and Model Stealing
@@
The outputs of a model, like the representation of textual data that a Transformer-based LM produces in the last layer, can be totally or partially inverted to reconstruct the text that is originally the input of the model. In particular, these representations – often called embeddings – are only apparently opaque and can be inverted to reconstruct the entire sentence or the words that it contains. Pan et al. (2020) design the first model inversion attack against Transformer-based Language Models. They observe that directly framing the inversion problem as an optimization problem is challenging due to the discrete nature of sentences. They train an attack classifier that is designed to extract fixed patterns and keywords – like, for example, location or medical information – from sentence representations of models. Similarly, Song & Raghunathan (2020) predict unordered sets of words from sentence embeddings: their attack model is a neural network trained with the multiset prediction loss. While the experiments in Pan et al. (2020) and Song & Raghunathan (2020) are performed on relatively smaller Transformer-based network compared to more recent models, their methods can generalize to larger networks with some additional difficulty due to the increasingly larger embedding dimension in LLMs.
@@
In some cases, the entire input text is also reconstructed. Li et al. (2023d) implement the inversion as a generative language modelling task with a decoder-only model, conditioned on the embedding of the sentence to invert as the first token representation. They show that the model successfully inverts meaningful tokens containing a large fraction of the named entity. The threatened model in their experiment is a BERT model, but the general framework can be extended given model representations for sentences. Morris et al. (2023) propose an iterative method to reconstruct the input text of Transformer-based models that produce embeddings for documents: their attack is based on the idea of iteratively generating different hypotheses that may justify the observed embeddings. They learn the distribution of texts in a dataset given their embeddings: their approach starts by assuming an initial hypothesis and then iteratively improves it by adjusting the hypothesis to progressively align its embedding to the target one. This approach has also been extended to a multilingual setting, adding a translation step before iterating again over the generated hypothesis Chen et al. (2024b). Only observing the logits of LLMs can also be sufficient to reconstruct the original input: Morris et al. (2024) train an inversion model on a collection of instruction datasets to reconstruct the prompt of an LLM – in their experiments the seven billion version of Llama-2 and Llama-2 Touvron et al. (2023) – given the probability distribution of the token following the prompt. In a completely black box scenario – in which only the model’s output is observed – an attacker may be able to reconstruct the original prompt training an inverter – another LM – on previously observed outputs on known input prompts Zhang et al. (2024a).
@@
This kind of attack has also been applied to Language Models fine-tuned on classification tasks to reconstruct private training data. In this case, the output of the model is a probability distribution over each of the possible classes. This information can be effectively exploited: in particular, an attacker may exploit the greater confidence that a model has in classifying sentences that are already part of the training set. Elmahdy et al. (2022) inject canaries in the training data (sequences that are intended to be later extracted and out of distribution with respect to the other data) and then reconstruct a partially masked sentence to search for tokens that maximize the probability of the target label. They perform the reconstruction by enumerating tokens in the vocabulary, weighted by the number of occurrences that they have in the training data. Zhang et al. (2022c) propose instead to build a dataset that resembles the unknown training data and initially train a text generator on these collected data, similarly to what is done with shadow model MIA (§3.2). Then, they further train the text generator via word embedding perturbation with the objective of minimizing the classification loss on the target model. The idea is that the target model’s prediction on training sentences will exhibit a lower loss, and this information can be used to perturb the text generator model and obtain data closer to the private one. Although the use of generative models as classifiers is not widespread at the time of writing, the results of this works could be generalized to the setting of generative LLMs with some additional difficulties in defining the target class of fine-tuned generative models.
@@ 3.3.2 Gradient Inversion
In Federated Learning (Ray et al., 2021), multiple nodes train a model cooperatively without directly sharing data samples: the system is designed to make training feasible while exchanging model updates rather than data across nodes (see §2.3 for further details). Gradient Inversion attacks are directed against federated learning environments that perform training by sharing gradients. Those attacks can be used to reconstruct training data from seemingly opaque sources.
@@
Zhu et al. (2019) demonstrated that if the gradients are openly accessible, it is possible to reconstruct the training data. The attacker aims to reconstruct the training data by minimizing the distance between some generated gradients and the ground truth one. In particular, the attacker initially generates gradients from random inputs and labels. Then, having access to the ground truth gradients solves the optimization problem of minimizing the $L_{2}$ distance between the generated gradients and the ground truth one, parameterized by real inputs and labels. In their experiments, Zhu et al. (2019) applied this attack against a BERT model. This idea was further explored by inverting gradients for image classification tasks (Zhu & Blaschko, 2021; Geiping et al., 2020; Yin et al., 2021).
@@
In reconstructing textual data, Deng et al. (2021) further attempted to perform a gradient attack against Transformer-based language models, introducing also a $L_{1}$ term in the reconstruction loss. Experiments on a Transformer model, BERT model and TinyBERT demonstrate the effectiveness of this gradient attack, recovering up to 50% of the original tokens on average. In LAMP (Balunovic et al., 2022), more effective results are achieved by simultaneously training the attack model to minimize the difference between the reconstruction gradients and choosing at each iteration only sequences that have low perplexity according to an external language model (like GPT-2). Reconstructing the correct word order is, in fact, one of the main challenges: Gupta et al. (2022) proposes to recover from the gradients a bag of words for the sentence to extract and then perform a beam search to effectively reconstruct the sentence.
Should be noted, however, that the moment of writing the application of those attacks to larger models is mostly limited by the increasing usage of centralized pretrained models.
3.3.3 Model Stealing
As exposing only black-box access to a LLM is a standard practice for popular commercial models, an adversary might want to reconstruct the gated model with the aim of imitating its behaviour. A family of attacks called model stealing attacks have been designed with this goal. The adversary, in this case, can have no information regarding the original training data. The attack usually requires observing the model outputs on inputs that the attacker fed to the gated model to reconstruct its internal parameters. Tramèr et al. (2016) introduce this family of attacks and demonstrate that an attack could lead to an accurate reproduction of logistic regression, neural networks, and decision trees. They also successfully demonstrate on those systems that model inversion as defined by Fredrikson et al. (2015) could be applied to the stolen model to reconstruct private input, with direct privacy implications.
Krishna et al. (2020) demonstrate the feasibility of those attacks against Transformer-based models. They use random words as input for gated fine-tuned BERT models and use their outputs for fine-tuning an equivalent model. They also demonstrate that the quality of the stolen model increases if the attacker has some background knowledge regarding the domain of the original fine-tuning dataset. Moreover, Zanella-Beguelin et al. (2021) notice that an attack that exploits the characteristics of the Transformer’s last layer can effectively reconstruct a fine-tuned BERT model with frozen pre-trained weights and learnable embedding projection matrix. In particular, they reconstruct the embedding projection matrix by feeding the gated model random inputs and collecting the corresponding output. They reconstruct the embedding projection matrix optimizing its weight on the observed embedding and outputs. Their technique, in combination with other techniques like the one proposed by Krishna et al. (2020), can lead to a refined model that resembles the target one.
Recently, a similar idea has been applied to threat LLMs: Carlini et al. (2024) proposed an attack to recover the embedding projection matrix at the last layer of a transformer language model. They attack a model in a black-box fashion, collecting the model’s output logits on random prompts. They notice that since the embedding projection matrix is low rank – since it projects the hidden dimension to a logit vector that has higher dimensionality – it’s possible to apply the SVD algorithm to the collected logits to obtain a linear transformation of the original embedding projection matrix. Moreover, they demonstrate that the logits over the entire vocabulary are not necessary, and that the probabilities over a predetermined number of words are sufficient for a successful attack. A similar approach has also been proposed by Finlayson et al. (2024), and they describe the utility of the attack in discovering whether a gated model has undergone some update.
3.4 Privacy Threats at Inference Time.
In the LLM context, the leakage of the original training data is not the only possible threat. The inclusion of additional, potentially private information at inference time presents a number of challenges. Here, we summarize three of the major challenges derived from such systems. These threats have pushed a great deal of research to focus on how to perform calculations securely on sensitive user input, either by obscuring them through differential privacy (§5.2) or by keeping them encrypted during the entire calculation process (§5.5).
Computation over Private Input.
Given the challenges of training such models from scratch, the option of using LLMs in a Machine Learning as a Service modality is becoming increasingly popular. In this setup, an organization hosts the model and users simply submit their data for processing. This approach raises significant privacy concerns for user data: vulnerabilities in the systems storing this data could lead to breaches (OpenAI, 2023), and model providers must offer strong assurances to comply with privacy regulations, such as GDPR. Similar concerns even led to a temporary ban on ChatGPT in Italy (Lomas, 2023).
This issue is particularly concerning for businesses dealing with proprietary data or commercially valuable information (Mitchell, 2023; Coles, 2023), and companies have decided, in some cases, to warn their employees about the responsible use of LLMs (Kim, 2023).
Prompting Private Data.
At inference time, the versatility of LLM is enhanced by the use of additional data, which are often provided in the form of demonstrations in the context of the task the user wants to perform (Brown et al., 2020). While being a useful technique, some sensitive information could be included in these examples in context as well, and an increasing body of research suggests that some risks can be identified during inference. In particular, since in-context learning is more cost-effective than fine-tuning, some systems may deploy LLMs that handle user requests by utilizing pre-loaded in-context demonstrations, which may include private or sensitive information. This efficacy, however, comes with an increased risk in terms of privacy leaks, which can also outweigh the risks against fine-tuned models (Duan et al., 2023). Wu et al. (2024b) propose a threat model in which a user of a system submits some data containing private information, such as medical records, and another user of the same system could potentially access that sensitive data. They focus on addressing this threat through the application of Differential Privacy during inference (see §5.2 for further details). Duan et al. (2024a) demonstrate the efficacy in a membership inference attack (MIA) against prompts containing sensitive, in-context learning demonstrations: they suppose that a model is prompted for a classification task and that the adversary may want to understand whether a certain example has been used or not as in context demonstration. They argue that an adversary can test whether an example has been used in context since the probability for the corresponding class will be much higher for members (in context examples) than for non-members. Similarly, in this scenario, a malicious user –instead of performing the intended task on the system– can potentially extract the in-context demonstrations via jailbreaking attacks (Tang et al., 2024). Chu et al. (2024) demonstrate that entire conversations can be reconstructed via jailbreaking when they are used as contextual information for prompting an LLM.
Contextual Privacy.
As LMs’ capabilities increase and their role as assistants becomes more prominent, it is important to understand whether they are able to distinguish if a piece of certain sensitive information is safe to be shared in a given context. For example, there is no privacy leakage when a doctor shares with his patients information about their own health. However, the disclosure of such information to an unauthorized insurance company violates patients’ privacy. Those considerations are recently emerging as language models can be reliably used as assistants by groups of users that want to share some information while preserving others. Brown et al. (2022b) discuss the discrepancy between data sanitization and differential privacy that assumes a structure of private data and cannot solve the necessity to share private information only with the correct group of people. They also argue that some data protection techniques – like data sanitization and differential privacy – make assumptions on the nature of the data that should be protected: implicitly or explicitly, they state that private information is structured information that should not be shared. However, every bit of language can potentially contain secrets that are not meant to be shared indiscriminately. Mireshghallah et al. (2024) formalize the necessity to share secrets in the appropriate context stemming from the contextual integrity theory (Nissenbaum, 2004). In particular, the information to be protected is not only in the training data but also during inference. They introduce a new benchmark to assess the ability of such models to disclose private information only to users who should be allowed to access them. They measure that multiple LLMs include information that should not be disclosed when asked to generate summaries of meeting conversations. Priyanshu et al. (2023) similarly explore the tendency of ChatGPT to disclose personally identifiable information when asked to produce summaries of hiring decisions and medical notes, a behaviour that they notice can be mitigated by explicitly asking the model to comply with privacy regulations and feeding step-by-step prompts that suggest which information should be deleted. Wang et al. (2024a) also discusses the necessity of measuring the privacy leakages during inference. They focus on PII like emails and phone numbers: they propose to initially feed the model with prompts that contain PII with the indication that those should not be shared; then, they prompt the model asking for the same information. They notice that GPT-3.5 is prone to violating privacy restrictions, while GPT-4 is more robust. Both models, however, tend to reveal private information when prompted with some demonstrations containing leakages in the few-shot scenario.
@@ 4 Solutions for Preserving Privacy in LLMs: Data
After identifying the key vulnerabilities and potential attacks on LLMs as outlined in §3, in this section, we introduce the reader to the current possible solutions to the problem of preserving privacy in the context of LLMs and text generative models. We have to say that the field is fragmented as this is a recent concern, and the field is evolving fast. At the moment of writing this survey, there is no single gold standard solution to the problem of preserving privacy in LLMs and text generative models. The main solutions available can be broadly categorized into two categories: i) Preserving Privacy in Data §4 and ii) Preserving Privacy in Model §5. In the former section, we collected all the methods that try to solve the problem at its source, i.e., they act directly on the privacy-sensitive information in the dataset before the training happens, while in the latter, we selected all the methods that address the problem from the model perspective, i.e., try to remove from the model problematic information after the training. The two approaches fit different use cases, so it is important to be aware of both.
In this section, we propose several solutions that allow for preserving privacy from the data perspective, i.e., we preserve privacy directly in the training data used in machine learning models. This includes any approach that involves manipulating, carefully choosing, or generating training data in order to avoid the memorization and subsequent leakage of private information inside a large language model or any model that learns from data. This basically puts the focus on preserving privacy in the data used to train the model. Given that a machine learning model is moulded by its training dataset, training it on a privacy-preserving dataset ensures that these methods offer equivalent privacy assurances to the dataset itself. The methods presented here are thus not limited to LLMs as they can be used to anonymize textual data more generally. That is, it is possible to use them to strip datasets of sensitive information before using these datasets to train LLMs. Making the training data safe means making the trained model safe.
In the context of LLMs, there are very few technical reports that thoroughly describe the precautions necessary to preserve privacy via data anonymization, and even fewer that release the preprocessing software used for this anonymization. Among the notable examples, the LLaMA3 technical report (Dubey et al., 2024) mentions "filters designed to remove data from websites [to eliminate] unsafe content or high volumes of PII" but does not detail the exact procedure while noting the use of "Microsoft Presidio Analyzer" in the preprocessing of speech data (See our overview of the tool in §6). Similarly, Soldaini et al. (2024), in the context of training OLMo (Groeneveld et al., 2024), also considered using Microsoft Presidio but ultimately opted for custom regular expressions, given the scale of the pre-training dataset.
@@
Overall, the available open reports suggest that for removing PII from the pre-training dataset, the preferred solution at scale tends to involve custom regular expressions (Penedo et al., 2024). The broader subject of preserving privacy in pre-training data, however, remains largely unexplored beyond these rudimentary techniques (Subramani et al., 2023). The scope of this section is thus two-fold: i) propose to the reader possible new avenues to preserve privacy in the pretraining dataset by surveying existing approaches proposed in the literature that have not been applied yet in the context of the pretraining phase of LLMs; ii) propose to the reader anonymization techniques that can be used to anonymize data for the instruction-tuning / fine-tuning phase where the LLMs are possibly fine-tuned on sensitive in-domain data for the downstream task (e.g., healthcare). Since most of these techniques were not tested with training datasets for LLMs, it is difficult to discuss about the scaling of these methods for big datasets, but seeing examples like the previously cited OLMo (Groeneveld et al., 2024), it is quite clear that complex methods at a big scale tend to not be chosen. However these techniques can be applied to data regardless of the subsequent usage, which means that they could (and will) be used to clean data for inference, before feeding it to a model.
The section is structured in the following way: we first show classical anonymization techniques §4.1, then we focus on anonymization techniques built with Differential Privacy §5. Putting this last section in relation with Figure 2, the point of application of DP would be (1), the input, because all these techniques aim to apply Differential Privacy on the data before feeding it to the LLM (for either training or inference).
Figure 4: Classical Anonymization Pipeline. The figure illustrates the standard process for anonymizing data in natural language processing tasks. The pipeline consists of three sequential blocks: Preprocessing, Identification, and Anonymization. The Preprocessing block prepares the raw data for further analysis. The Identification block involves the use of a procedure to detect potential Personally Identifiable Information (PII) within the data (e.g., Named Entity Recognition). Finally, the anonymization block replaces identified PII with anonymous substitutes, ensuring data privacy preservation. Arrows indicate the flow of data through each stage of the process.